Please help with problem:
Lisa has 9 rings in her jewelry box. Five are gold and four are silver. If she randomly selects 3 rings to wear to a pary, find each probability:
a) P(2 silver or 2 gold)
b) P( at least 1 silver)

Answer by stanbon(75887)

Please help with problem:
Lisa has 9 rings in her jewelry box. Five are gold and four are silver. If she randomly selects 3 rings to wear to a pary, find each probability:
a) P(2 silver or 2 gold)
# of ways to succeed: 5C2*4C1 + 4C2*5C1 = 10*4 + 6*5 = 70
# of possible outcomes: 9C3 = 84
P(2 silver or 2 gold) = 70/84
-----------------------------------
b) P( at least 1 silver) = 1 - P(no silver) = 1 - [5C3/9C3] = 1-(10/84)
= 74/84 = 37/42
====================
Cheers,
Stan H.
======================